# Situationally-Aware Dynamics Learning

### v2

@@ VI Discussion
Limitations. While the proposed approach demonstrates significant advantages, it is not without its limitations. In simulation, we observed a slight decline in path smoothness metrics, even as the overall success rate remained high. This trade-off appears to stem from the situation changes, which trigger re-planning with a new set of constraints provided by the dynamics model associated to the updated situation. These transitions, while enabling emergent adaptive behaviors, occasionally resulted in less smooth trajectories. However, this effect did not critically impact the robot’s performance or overall mission success, suggesting that the benefits of adaptability outweighed the costs associated with these deviations.
VII Conclusion
@@
VIII Acknowledgement
The authors acknowledge Youwei Yu for his support during the real-world experiments.
IX Declaration of Conflicting Interests
@@
X Funding
This work was supported by NSF: CAREER: Autonomous Live Sketching of Dynamic Environments by Exploiting Spatiotemporal Variations (grant #2047169). This work used Amazon Web Services through the CloudBank project, which is supported by National Science Foundation grant #1925001.
Appendix A Additional Experimental Results
@@
A-A Lake Simulator
Defining and identifying the hidden state in the unstructured terrain navigation task is inherently challenging due to the complexity of natural environments and their dynamic properties. Therefore, to further validate our online hidden state representation learning method, we conduct experiments in a controlled environment where we can directly affect the UDGP of the world. For this, we use the Lake World Simulator [bingham19toward], as shown in Figure 17. This environment provides a practical testbed for analyzing navigation in maritime settings, where dynamic and unpredictable conditions are prevalent. Specifically, we simulate a scenario in which an Autonomous Surface Vehicle (ASV) must navigate across the lake while responding to shifting wave patterns about which it does not have any kind of information beyond what it can measure via its proprioceptive odometry measurements. The ASV’s objective is to navigate toward randomly generated targets while managing to do so with the changing wave patterns.
The wave generation patterns in this simulation are governed by parameters that are intentionally randomized during the experiment to induce shifts in the UDGP of the transition dynamics. These shifts represent transitions between different hidden states, such as calm waters, mild ripples, or high waves, which directly impact the ASV’s transition dynamics. Our method needs to obtain hidden state representations that encode this circumstances in real time. This experiment is particularly meaningful because it reflects real-world challenges faced in maritime robotics, such as navigating through varying sea states where external conditions like wind, currents, and waves cannot always be directly measured.
@@ A-A Lake Simulator
A-B Difficult Terrain Navigation
Figure 7 in the main manuscript illustrates changes in detection and modeling results. To preserve the manuscript’s flow and address space limitations, Figure 7 presents only the first six dimensions of the 26-dimensional observations used for online situation modeling and identification. Figure 19, however, presents the full observation history, offering a more detailed view that highlights how effectively our multivariate extension of BOCD detects changes in the UDGP of the robot’s dynamics during an unstructured terrain navigation task.

### v3

@@ 7.1 Limitations and Scope
While the proposed approach demonstrates significant advantages, it is not without its limitations. In simulation, we observed a slight decline in path smoothness metrics for the UGV evaluation, even as the overall success rate remained high. This trade-off appears to stem from the situation changes, which trigger re-planning with a new set of constraints provided by the dynamics model associated to the updated situation. These transitions, while enabling emergent adaptive behaviors, occasionally resulted in less smooth trajectories. However, this effect did not critically impact the robot’s performance or overall mission success, suggesting that the benefits of adaptability outweighed the costs associated with these deviations.
A more fundamental limitation relates to the model’s generalization capabilities, particularly when encountering dynamics that are out-of-distribution (OOD) with respect to the situations learned during training, which is a common restriction of most learning-based methods. Our SA framework excels at identifying and representing latent factors that are either represented within its context library or can be inferred from online transition data. However, its performance will predictably degrade if the test-time dynamics are entirely outside the support of this learned distribution. Such scenarios could include encountering fundamentally different physical phenomena (e.g., transitioning from rigid-body to deformable contact mechanics), experiencing novel forms of actuator or sensor degradation not seen during training, or operating in an environment with entirely new external forces. However, when the robot encounters a new environment, its observed state transitions are associated with the most probable situation symbol from the existing library. The planner then utilizes a dynamics model conditioned on this best-match context/situation. For instance, Figures 3 & 21 illustrate that the dynamics for tiles and concrete are closely related, while both are distinct from those of pebbles. Consequently, if a robot trained on concrete were to encounter tiles for the first time, the system is expected to reuse the same situation it employed for ‘concrete’. This mechanism is intended to reduce compounding error in the dynamics model and improve downstream planning performance.
In practice, OOD conditions can be diagnosed by monitoring key statistical indicators, such as a sustained increase in the predictive uncertainty of the dynamics model or a significant divergence between observed state transitions and those predicted by any available context. To mitigate this, the framework is amenable to established techniques; for instance, the context library can be made more robust by proactively expanding its support during training using methods like domain randomization to cover a wider array of potential dynamics. For truly novel situations, this diagnostics could trigger a safe-mode or a request for human intervention, forming a clear direction for future work in lifelong adaptation.
8 Conclusion
@@
9 Acknowledgement
The authors acknowledge Youwei Yu for his support during the real world experiments.
10 Declaration of Conflicting Interests
@@
11 Funding
This work was primarily supported by NSF: CAREER: Autonomous Live Sketching of Dynamic Environments by Exploiting Spatiotemporal Variations (grant #2047169) and partially supported by Army Research Lab (grant #W911NF2520083). This work used Amazon Web Services through the CloudBank project, which is supported by National Science Foundation grant #1925001.
Appendix A Latent Factor Effect Characterization
Here we seek to present a real world example of the motivation for a latent factor representation learning method. Specifically, we replicated the motivating example from Figure 1 with a UGV traversing flat terrain. There we explained that executing the same action from similar states can yield significantly different outcomes due to the presence of latent, unmodeled factors (e.g., wind, terrain friction, actuator wear). Our method shows that by explicitly representing these latent factors and additionally conditioning the dynamics model on that representation helps reduce the prediction error and improve the control performance in unstructured environments.
We proceed to describe our focused real-world experiment that directly illustrates the phenomenon and motivates the conditional modeling approach: we collected 250 state transitions from comparable states after applying the same action ($v_{x}=0.5\ \mathrm{m/s}$, $\omega_{z}=0.0\ \mathrm{rad/s}$) across 14 different environments, for a total of $3500$ state transitions. These environments are presented in Figure 20. The Supplementary Video 4 (SV4) showcases this data collection process: https://youtu.be/xGSFT9QvtJ0. This video illustrates our best effort at a controlled data collection process for a single $(s,a,s^{\prime})$ transition, where terrain type serves as the latent factor. The goal is to demonstrate that even for this simple case, aggregating all collected data without accounting for the latent factor can significantly affect predictions. Overall, this let us see that conditioning the dynamics model on the latent factor representation yields predictive improvements. While these gains are expected to be even more pronounced in challenging scenarios, this simple setup highlights that the issue already arises in the most basic case.
Figure 21 visualizes the resulting outcome distributions for this single case, to show where distributions align and where they diverge, seeking to explicitly demonstrate the effect of the latent factors after applying the same action from comparable states [^3] We cannot say they are the same state as there are natural differences, but we tried to replicate them as closely as possible. Also, as discussed in Section 5.3, the state is preprocessed to make it position invariant.. The figures demonstrate that, while some environment pairs produce similar outcomes, others diverge substantially — exactly the behavior that motivates conditioning on an inferred situational representation. In particular, if we were to approximate the dynamics for this particular case using all the collected data (without conditioning on latent factors) as is traditionally done, it would fail to accurately model all dynamics. Even though this might not appear like a critical problem in this isolated motivating example, our overall results have shown that the problem is exacerbated in more complex domains and platforms, such as uneven terrain navigation or flying under evolving wind conditions.
Figure 3 also demonstrates this effect. Colored curves correspond to different terrain types, and the dashed black curve denotes the unconditioned (aggregated) estimate, demonstrating how aggregation obscures terrain-specific transition dynamics. Note that we do not claim that every terrain type requires a separate latent-factor representation. Instead, there is a trade-off between the number of modeled latent factors and their predictive benefit. For example, Figure 21 shows that the distributions for tiles and concrete are very similar, suggesting that a single representation may be sufficient in this case. Crucially, such discrepancies are often better attributed to an incomplete state representation—e.g., the robot lacks sensors, such as a camera, to detect terrain types—rather than inherent aleatoric errors. Additionally, note that our approach does account for aleatoric uncertainty in the transition model, as done in traditional dynamics learning methods; however, our primary focus is on reducing epistemic uncertainty by identifying and accounting for the latent factors influencing the dynamics.
Figure 20: Environments used for data collection to study the influence of latent factors (in this case, terrain type) on UGV dynamics. In each environment, the same action was executed for 50 timesteps starting from a comparable initial state, and the procedure was repeated five times per environment.
Figure 21: Descriptive statistics of the effect of the latent factor on the dynamics of the UGV. Data obtained from 250 state transitions after applying the same action $v_{x}=0.5\penalty 10000\ m/s$ and $\omega_{z}=0.0\penalty 10000\ rad/s$.
Appendix B Additional Experimental Results
B.1 Lake Simulator
@@
Defining and identifying the hidden state in the unstructured terrain navigation task is inherently challenging due to the complexity of natural environments and their dynamic properties. Therefore, to further validate our online hidden state representation learning method, we conduct experiments in a controlled environment where we can directly affect the UDGP of the world. For this, we use the Lake World Simulator Bingham et al. (2019), as shown in Figure 22. This environment provides a practical testbed for analyzing navigation in aquatic settings, where dynamic and unpredictable conditions are prevalent. Specifically, we simulate a scenario in which an Autonomous Surface Vehicle (ASV) must navigate across the lake while responding to shifting wave patterns about which it does not have any kind of information beyond what it can measure via its proprioceptive odometry measurements. The ASV’s objective is to navigate toward randomly generated targets while managing to do so with the changing wave patterns.
The wave generation patterns in this simulation are governed by parameters that are intentionally randomized during the experiment to induce shifts in the UDGP of the transition dynamics. These shifts represent transitions between different hidden states, such as calm waters, mild ripples, or high waves, which directly impact the ASV’s transition dynamics. Our method needs to obtain hidden state representations that encode this circumstances in real time. This experiment is particularly meaningful because it reflects real-world challenges faced in aquatic robotics, such as navigating through varying sea states where external conditions like wind, currents, and waves cannot always be directly measured.
@@
B.2 Difficult Terrain Navigation
Figure 8 in the main manuscript illustrates changes in detection and modeling results for the UGV. To preserve the manuscript’s flow and address space limitations, Figure 8 presents only the first six dimensions of the 26-dimensional observations used for online situation modeling and identification. Figure 24, however, presents the full observation history, offering a more detailed view that highlights how effectively our multivariate extension of BOCD detects changes in the UDGP of the robot’s dynamics during an unstructured terrain navigation task.